OMG Inc. has 6 million shares of common stock outstanding, 5 million shares of preferred stock outstanding, and 7,000 bonds. Suppose the common shares are selling for $17 per share, the preferred shares are selling for $26 per share, and the bonds are selling for 108 percent of par.
What would be the weight used for equity in the computation of OMG’s WACC? (Round your answer to 2 decimal places.)
Weight used: ____.__%
Answer #1

Total value of:

Common stock=(6*17)=$102 million

Preferred stock=(5*26)=$130 million

Bonds=7000*(1000*108)%=$7.56 million

Total value=(102+130+7.56)=239.56 million

Hence weight of equity=$102/239.56

=42.58%(Approx).